Two jets pass directly over each other at 10 pm. One jet travels south at 480 km/hr, the other travels east at 900 km/hr. How far apart are the two jets at midnight?
Answer by rfer(16322)

480*2=960
900*2=1800
1800^2+960^2=c^2
3240000+921600=4161600
c=sq rt 4161600
c=2040 km